### v2

# Curating art exhibitions using machine learning

Abstract
Here we present a series of artificial models – a total of four related models – based on machine learning techniques that attempt to learn from existing exhibitions which have been curated by human experts, in order to be able to do similar curatorship work. Out of our four artificial intelligence models, three achieve a reasonable ability at imitating these various curators responsible for all those exhibitions, with various degrees of precision and curatorial coherence. In particular, we can conclude two key insights: first, that there is sufficient information in these exhibitions to construct an artificial intelligence model that replicates past exhibitions with an accuracy well above random choices; and second, that using feature engineering and carefully designing the architecture of modest size models can make them almost as good as those using the so-called large language models such as GPT in a brute force approach.
Keywords: art curation; machine learning; embeddings; large language models (LLMs)
1 Introduction
Artificial intelligence (AI) and its applications are currently a hot topic in academia as well as in the news, particularly with the announcement of the 2024 Nobel Prize in Physics, which was attributed to work by researchers in AI [The Royal Swedish Academy of Sciences, 2024]. In the area of art, AI has already been applied in several research fields, including: generating art (images) in any style one wants [Gatys et al., 2016]; generating images from pure text [Ramesh et al., 2021, Esser et al., 2021, Radford et al., 2021]; translating paintings to possible corresponding real-world images (or vice versa) and other artistic styles [Zhu et al., 2017]; discovering similar patterns in disparate paintings [Saleh et al., 2016, Shen et al., 2019]; clustering similar artworks [Hamilton et al., 2021, Castellano and Vessio, 2022]; style/artistic school and artist identification/classification [Keren, 2002, Shamir et al., 2010, Shamir and Tarakhovsky, 2012, Cetinic and Grgic, 2013, Saleh and Elgammal, 2015, The Museum of Modern Art, 2025b]; painting to text/semantic knowledge extraction [Vinyals et al., 2015]; aesthetic evaluation and artistic opinion extraction [Cetinic et al., 2019]; and museum/exhibition curation and artwork selection/recommendation [He et al., 2016, Krysa and Moscoso, 2019, Messina et al., 2019, 2020, Bowen et al., 2020, Fosset et al., 2022, Yilma and Leiva, 2023, Nasher Museum of Art at Duke University, 2023, Schaerf et al., 2023, Herman and Moruzzi, 2024, von Davier et al., 2024, Srinivasan, 2024, Gordon, 2025, Li, 2025].
Regarding the last field of research, museum and exhibition curation using AI, most of the work seems to focus on either finding similar artworks to a given (seed) artwork, or using a prompt (e.g., for the OpenAI GPT-4 [OpenAI, 2023]), to select appropriate artworks. Here we propose a different and modest approach that does not seem to have been explored yet. We propose to use existing collections of actual past art exhibitions [see e.g., ArtLAS Project, 2025, University of Vienna, 2025, Museum of Modern Art (2025), MoMA, National Gallery of Art, 2025, The Metropolitan Museum of Art, 2025a, and others], to create a recommender system [for an example of recommender systems see Hu et al., 2008] using AI that imitates what human curators do. That is, we propose to teach AI to learn from past curatorship. The closest of the above-mentioned works to our proposal is the one by the Nasher Museum of Art at Duke University [Nasher Museum of Art at Duke University, 2023], which used OpenAI’s ChatGPT platform to curate exhibitions. The main difference with respect to our work here is that our modelling is absolutely restricted to a given set of artworks (therefore no hallucinations giving artworks outside that set), and more importantly and crucially, we have made the models learn from past exhibitions using those artworks and only learn from that set of information.
This is how we propose to train our models. The training set would consist of the input pairs $(x,y)$, where $x$ is the embedding vector [see e.g., Kenton and Toutanova, 2019, on text-to-vector embeddings] of the exhibition title and main text, while $y$ is made of the features extracted from the images of the artworks exhibited. Optionally, one could add to $y$ the embedding vector of the caption of artworks that were selected by the human curators, if available. [^1] We could consider also using the information on who was the human curator, for a more nuanced system, that could recognise and create an exhibition on the style of a certain human curator, but given the scarcity of information available, that is wildly optimistic. After the training phase, users could input on the test phase, a title and a description of a hypothetical exhibition, and the system would choose the appropriate artworks to include, and would choose only from the given set of artworks.
2 Dataset
Although there are several datasets available [see e.g., ArtLAS Project, 2025, University of Vienna, 2025, Museum of Modern Art (2025), MoMA, National Gallery of Art, 2025, The Metropolitan Museum of Art, 2025a, and others], we have decided to restrict ourselves to the Metropolitan Museum of Art in New York (hereafter Met Museum), as it had two essential ingredients for our intended research project, namely: an extensive and well formatted dataset on past exhibitions and their works of art, and the corresponding freely available dataset on details and tags associated with those works of art, i.e., their metadata. Furthermore, the merging of databases could have proven a very complex task, as these sets all have different metadata associated to the artworks, and the data formats come in a large variety.
On the first part of this dataset, we have focused on the data from the Met Museum’s website [The Metropolitan Museum of Art, 2025b], and extracted directly the full set of archived exhibitions and where available, its corresponding works of art (or rather the object id, an integer that clearly identifies the artwork within the Met Museum’s associated artwork database). This dataset included a total of $1636$ exhibitions from 2000 to 2025 inclusive, although these included duplicated exhibitions, as some of those encompass several years. Also, some exhibitions did not contain the list of object ids to map to works of art and/or only contained objects temporarily loaned to the Met Museum and therefore not in the database of objects. After removing those and other exhibitions we could not use, we end up with a much lower value of $236$ exhibitions. We depict in Table 1 a summary of that data, which corresponds to a download on the date of 24 February 2025 (naturally, the Met Museum may have released more webpages with newer exhibitions, but we have frozen our data on this date).
Year Total Exhibitions Exhibitions with Object IDs Works of Art Found Word Count
2000 55 6 $168$ $5911$
2001 59 11 $403$ $6554$
2002 54 11 $254$ $6242$
2003 56 9 $371$ $5911$
2004 57 6 $342$ $6052$
2005 49 4 $239$ $5544$
2006 49 5 $308$ $6147$
2007 51 16 $847$ $6019$
2008 57 19 $1203$ $6947$
2009 61 18 $863$ $6905$
2010 61 16 $1307$ $6246$
2011 68 24 $1574$ $7246$
2012 65 30 $2661$ $7373$
2013 79 39 $2863$ $9301$
2014 94 53 $2827$ $9832$
2015 90 53 $2741$ $8663$
2016 95 58 $3646$ $9103$
2017 95 2 $111$ $9694$
2018 78 2 $129$ $8222$
2019 74 2 $129$ $7462$
2020 56 5 $276$ $5876$
2021 58 26 $2042$ $9727$
2022 56 47 $3711$ $12,796$
2023 60 47 $3483$ $14,078$
2024 50 36 $2986$ $10,434$
2025 9 8 $736$ $1590$
Total 1636 553 $36,220$ $199,875$
Total after removal of multi-year exhibitions 1009 338 $20,172$ $123,349$
Total after removal of artworks not on Met Museum 236 $10,470$ $26,388$
Table 1: Exhibition statistics per year for the downloaded exhibitions from the Metropolitan Museum of Art of New York, which corresponds to a download on the date of 24 February 2025. It shows the total number of exhibitions per year, the number of those that have the artworks explicitly stated (using Object IDs that match the artwork database from the Met Museum) and the number of artworks found. As some exhibitions run across year ends, we also show the corresponding total all years’ value of the number of exhibitions that have known Object IDs and the corresponding lower number of actual artworks in that unique set of exhibitions. Furthermore, we also show those numbers when one removes exhibitions that encompass only artworks not on the Met Museum’s database, i.e., exhibitions that are made of loaned artworks. The actual usable number of exhibitions for our research project is therefore only 236, and the number of artworks a total of $10,470$ (with possible repeats across exhibitions). The number of actual unique artworks is $9289$. Finally, we show the number of words in each year for all the titles and descriptions for all exhibitions within that year, and the total number of words for all exhibitions with and without duplication, and for the actual exhibitions we shall use for our modelling. For comparison, it is speculated that the OpenAI’s gpt-4o-mini model uses several billion words for its training.
The exhibition dataset, consisting of a grand total of 236 exhibitions after removal of multi-year duplicates and made of exhibitions for which we have object ids in the corresponding Met Museum’s artworks object id to metadata file, has the following format, as depicted in Listing 1. For each exhibition, we have the title of the exhibition and the description of the exhibition (called overview_text), all in English and with the replacement of any diacritics (accents, umlauts, tildes, etc.) and non-ASCII characters, and then we have a variable number of object_ids, which identify the artworks within the Met Museum. [^2] Note that this id is not a worldwide id, but an internal Met Museum id. There are some attempts to catalogue and identify all artworks with a unique worldwide id, e.g.,: Wikidata Qxxxxx ids — “Les Demoiselles d’Avignon” by Pablo Picasso has a unique id Q910199; the International Image Interoperability Framework [IIIF Consortium, 2024]; the Europeana project [Europeana Foundation, 2024]; among many others. For each object_id, we have a series of metadata elements, which can be a single string, a list of strings or an empty list with no elements in it.
"exhibitions": [
"title": "Sculpture and Decorative Arts of the Spanish Renaissance",
"overview_text": "The Metropolitan Museum of Arts small but excellent collection of Spanish polychrome sculpture, including sacred reliefs and freestanding carved figures once housed in the churches of Spain, is displayed in the gallery adjacent to the newly reopened Velez Blanco Patio. The selection, which displays the unique blending of early western European and Islamic stylistic and technical influences, emphasizes the diversity in the material culture of Renaissance Spain after the Catholic reconquest by Ferdinand and Isabella.",
"object_ids": {
"187702": {
"Department": "European Sculpture and Decorative Arts",
"Object Name": "Jug",
"Title": "Jug",
"Artist Display Name": [],
"Object Begin Date": "1600",
"Medium": "Tin-glazed earthenware",
"Classification": [
"Ceramics-Faience"
"Tags": [
"Cranes",
"Donkeys",
"Trees"
"187863": {
"Department": "European Sculpture and Decorative Arts",
"Object Name": "Bottle",
"Title": "Bottle (Refredador)",
"Artist Display Name": [],
"Object Begin Date": "1500",
"Medium": "Tin-glazed and luster-painted earthenware",
"Classification": [
"Ceramics-Pottery"
"Tags": [
"Coat of Arms"
Listing 1: The excerpt of the raw format of our data, showing an exhibition by the Met Museum, from the year 2000 (https://www.metmuseum.org/exhibitions/listings/2000/spanish-renaissance), displaying the title and overview text describing the purpose and essence of the exhibition, and some of the mappings to the actual object_ids and their respective metadata from The Metropolitan Museum of Art [2025a].
On the second part of this dataset, we used the Met Museum’s Open Access CSV in The Metropolitan Museum of Art [2025a] data, which contains the museum’s entire collection index of works of art (a total of $484,956$ works of art) and its associated metadata. The majority of these objects did not show up on any exhibition on the list of exhibitions that we describe above. The metadata consists of many fields (54 in total), but we focus only on “Department”, “Object Name”, “Title”, “Artist Display Name”, “Object Begin Date”, “Medium”, “Classification”, “Tags”, together with “Object ID” to link to the exhibitions’ extract. Note that later on, for the machine learning approaches we ignore “Object Name”, “Title”, as these are somehow too arbitrary to be used as meaningful metadata for our modelling. The count of the top strings in each of the fields we consider (“Department”, “Artist Display Name”, “Object Begin Date”, “Medium”, “Classification”, “Tags”) is depicted in Table 2. We shall call these set of all these fields generalised tags (as opposed to only tags, a name which is already used by the Met Museum as a column name in its dataset of all artworks.) There are a total of $69,566$ generalised tag strings in our full dataset of 236 exhibitions, of which $8591$ are unique generalised tags strings. Below we display a very small subset of this dataset in Table 4 as an example for the reader.
Field Value Count
Department Asian Art 2893
Photographs 1720
Drawings and Prints 1634
European Sculpture and Decorative Arts 1506
Modern and Contemporary Art 725
Artist Display Name Pablo Picasso 484
Galerie Louise Leiris 273
Hidalgo Arnera 165
Aldo and Piero Crommelynck 126
American Tobacco Company 121
Object Begin Date 1700 617
1750 351
1800 347
1900 343
1600 199
Medium Gelatin silver print 449
Oil on canvas 277
Photomechanical print 250
Albumen silver print from glass negative 223
Etching 197
Classification Prints 1776
Photographs 1438
Paintings 1319
Drawings 660
Creche 464
Tags Men 2340
Women 1382
Portraits 797
Flowers 521
Landscapes 369
Table 2: Top 5 values per field or generalised tag count in the Met Museum exhibition data.
Field Non-empty Entries
Department $484,956$
Object Name $484,956$
Title $484,956$
Artist Display Name $282,513$
Object Begin Date $484,956$
Medium $484,950$
Classification $406,239$
Tags $192,455$
All Fields Non-empty $120,713$
Table 3: Non-empty Fields in Metropolitan Museum of Art Dataset. Notice that the fields “Object Name” and “Title” are used for information purposes only, and do not enter the machine learning approaches.
Generalised tags
Object ID 358770 483355 490021
Department Drawings and Prints Modern and Contemporary Art Modern and Contemporary Art
Object Name Print Painting Painting
Title Les Deux Baigneuses… Erotic Scene (La Douceur) Still Life with Mandolin and…
Artist Display Name Auguste Renoir Pablo Picasso Pablo Picasso
Object Begin Date 1895 1903 1924
Medium Etching on wove paper… Oil on canvas Oil and sand on canvas
Classification Prints Paintings Paintings
Tags Female Nudes|Bathing Men|Female Nudes Food|Mandolins|Still Life
Table 4: Excerpt of the Met Museum database of all its $484,956$ artworks. We select only a few metadata fields, which we call generalised tags. The total number of fields present on the database is 54 in total, but we use only “Department”, “Artist Display Name”, “Object Begin Date”, “Medium”, “Classification”, “Tags”, together with “Object ID” to link to the exhibitions’ extract. Above we display “Object Name”, and “Title” for information purposes only, these are not used in the modelling. We truncate some of the values of the metadata fields with “…” for presentation only in this table, we use the full strings in our modelling code. For some of the values we have more than one string, e.g., “Female Nudes|Bathing”, which means we have a list of strings, two strings in this particular case. The “|” or pipe symbol is used in comma separated files (csv files) as the one given by the Met Museum in its Application Programming Interface (API) [The Metropolitan Museum of Art, 2025a] to expose its database to external users.
The dataset is very extensive, and we are not interested in all of those fields, what we shall call hereafter generalised tags. In fact, not all generalised tags are filled for all artworks, which is not surprising, given that some artworks are centuries or millennia old. We summarise in Table 3 some of these statistics.
Our goal is to map, using machine learning techniques, the title and overview text, to the fields named above or what we call generalised tags, and therefore via those to select artworks from the full set of $484,956$ works of art in the Met Museum.
3 Methodologies and results
We attempt this mapping between inputs and outputs via four distinct approaches. All our approaches draw on methods developed in the field of Natural Language Processing (NLP) – see e.g., Young et al. [2018] for a review and references therein. Schematically, all our approaches can be summarised by Figure 1, which displays a toy model where we input an exhibition title and description, and output either a list of generalised tags (which is then mapped to the nearest objects that match those tags), or an actual direct list of artworks. Naturally, the neural networks we use are much larger, both in depth and width. Furthermore, some of those neural networks can be much more complex, incorporating other types of connections including loops.
Input Layer ($x$) Hidden Layer 1 Hidden Layer 2 Output Layer ($y$) Exhibition title and description List of generalised tags and/or probability of generalised tags and/or list of artworks
Figure 1: Example of a toy model neural network mapping exhibition data, in this case the title and description (overview text) of the exhibition to the generalised tags or a list of artwork suggestions to assign to the input exhibition. Each node or neuron on the neural network represents a real number, the level of the neuron activity, and the links between the nodes are called weights and represent the influence of that neuron on other neurons.
First, we use just the data as it is, without external help, by using its internal statistics. That is to say, we use the probability of finding a word in English within the corpus [^3] A corpus is a collection of texts that are used for linguistic/statistical analysis. of exhibition titles and descriptions. In other words, we create a model to represent those words within the language patterns of that corpus. In NLP, this is usually done by using the so-called text vectorisation, which computes word (more accurately, it uses tokens, which can be words, subsets of words or letters) frequency and relative distribution within the corpus [Salton et al., 1975]. Therefore, we create a text vectorisation based on the contents of all the titles concatenated with the overview (description) texts for all exhibitions taken together. That is to say, we transform the input texts into numeric vectors using the self-contained statistics of the exhibition titles and texts [see e.g., Mikolov et al., 2013, on embeddings]. Second, we now use external help, in the form of the state-of-the-art OpenAI embedding model text-embedding-3-large that can perform very high-quality text vectorisation. In this sense, we are using the statistical analysis performed by OpenAI on their (much) larger corpus of texts to determine the best (numeric) vectors that represent strings of text. These models also have the advantage of capturing semantic similarity in a metric way [see e.g., Corley and Mihalcea, 2005, for details of semantic similarity in NLP], i.e., they convert textual data into numeric vectors where similar words/sentences/paragraphs have similar vector representation, being near, in the high-dimensional space [^4] We note here that the dimension of these spaces can be very large, typically of the order of thousands or even tens of thousands of dimensions, making it impossible to clearly visualise these vector representations. characterising the texts. In our first model, we do not capture semantic similarity, only word frequency, so this second approach, has, in principle, some advantages. Third, we use OpenAI embedding models to transform both the input, the title and description of the exhibitions, and the output, for which we use the concatenated strings made of all the metadata associated with all the artworks for each of one the exhibitions. We then can find the nearest embedding vector for an input text and from that find the nearest artworks’ metadata embedding vectors for that text, giving us the nearest artworks. Fourth, and finally, we use the full power and knowledge encapsulated in the OpenAI models, by using the OpenAI Application Programming Interface (API [^5] An API is a collection or set of rules and specifications that allows to call code or databases remotely without having to store or host the entire infrastructure locally.) to call OpenAI models to ask for a direct and full list of generalised tags suggested for a given exhibition description (title concatenated with the overview text). We therefore train one of the latest OpenAI models, “gpt-4o-mini”, using a technique called fine-tuning [Devlin et al., 2019, Yosinski et al., 2014]. Fine-tuning models is a common approach in the field of neural networks, taking an already pre-trained network and improving it (or fine-tuning it) to a more specific domain [see e.g., Covas, 2020, 2023, for two examples in astrophysics and in finance]. After fine-tuning, we ask the resulting model via the API for lists of generalised tags (without forcing the number of predicted artworks) given an input text (which the fine-tuned model will decided how to represent and embed numerically).
We note that, for the first, second and third models, we do not take the approach of directly attempting to predict the full list of artworks from the input text, i.e., the text as depicted in Listing 1, or in other words, the number of artworks. This is because, first, it is very difficult to predict a variable number of text strings, i.e., the metadata for the several artworks. Actually, we found that there was no significant correlation between the statistics of the word distribution within the input texts and the number of artworks in each exhibition. Second, because the entire database of artworks is much larger ($484,956$ artworks than the number of artworks showing up in our exhibition dataset ($10,470$ artworks). Therefore, for the first, second and third models, we impose the number of artworks to be either the number of the artworks of the actual exhibition (for in-sample prediction [^6] This is what a prediction made using the same data that was used to train the model is called.) or 16 artworks (for out-of-sample prediction [^7] Conversely, this is what a prediction made on new data not used during model training is called. These kinds of predictions are made to assess model generalization.). For the fourth model, the full GPT model, we can let the algorithm decide itself.
For all approaches, we divide our dataset in Table 1 into a training and validation sets, as it is standard in machine learning approaches. We do this in the proportion $80\%$ examples for training (a total of 188 examples) and $20\%$ samples for validation or in-sample testing (a total of 48 examples), taking the total to the above-mentioned global set of 236 exhibitions. [^8] The validation set is a subset of examples drawn from the same sample – this means it presumably follows the same probability distribution function – but is not used for training, just for checking accuracy. On the other hand, test sets can be in-sample (from the same probability distribution) or out-of-sample (from possibly a different distribution), and the latter are used to test the ability of the model to actually generalise. We train the models on the training set, [^9] We train and optimise the neural network using the so-called Adam optimiser [Kingma, 2014], which is implemented within Google’s TensorFlow deep learning framework [Abadi et al., 2016] in our Python code. and then test our ability to predict the appropriate metadata (“Department”,“Artist Display Name”, “Object Begin Date”, “Medium”, “Classification”, “Tags”) and ultimately the artworks for the input text in the validation set, comparing the predicted versus actual metadata or generalised tags and the list of predicted versus actual artworks. All our training optimises to minimise the error difference between the actual output (on the training set) and the predicted output. We also calculate the error difference within the validation set to monitor over-fitting, a common problem with machine learning models [Geman et al., 1992]. Only after training the model, and checking the results for the training and validation sets, we then test out-of-sample on a made-up test set, where we create an input hypothetical exhibition title and description, in the hope that our model can select appropriate artworks for it.
3.1 Self-contained text vectorisation
Text vectorisation transforms human language in the form of text into a series of numbers, of fixed or variable length, that represent that text in a machine-readable format [Mikolov et al., 2013]. That way, we can transform the title and the overview text (concatenated together) into a numerical vector – a series of ordered numbers (and we choose this to be of fixed size to keep the model complexity under control and to have a fixed number of input nodes on our neural network). This will serve as the $x$ input in our machine learning approach, a neural network, as depicted in the conceptual toy model in Figure 1. As implied by the toy model diagram, we therefore also need an output $y$ variable. One approach would be to have the actual artworks and metadata to be a running text output, like what we have on the raw data as depicted in the Listing 1 above. However, this would introduce two hard-to-solve challenges: first, the output is of variable size (length), since an exhibition can have any number of associated artworks, and also the artworks can have any number of metadata items associated with them; and second it would not be possible to constrain the neural network to select only artworks from the available $484,956$ works of art in the Met Museum. The first issue could potentially be solved by a so-called Long Short-Term Memory (LSTM) neural network [Hochreiter and Schmidhuber, 1997]. These are used, for example, in one human language to another human language translation (or computer programming language to another), and have the advantage of allowing for variable input/output length. [^10] This ability to achieve variable output length is reached by using a neural network architecture with feedback loops, called a recurrent architecture [Elman, 1990]. The neural network outputs a token (or word) at a time, and then that output token is concatenated with the input and fed back into the network until a special end-of-sequence (EOS) token is produced. This allows the model to have a flexible output length. Nonetheless, even with a LSTM network, and with only a small number of training examples like we have, it would be very difficult to get it to output a consistent format like the one we have for our raw data, with all those fields and fixed format metadata. Furthermore, we really want to constrain our machine learning method to the range of Met Museum artworks only, and not to a hypothetical range of all world artworks. Therefore, we have devised a simple approach that consists of two steps, first in flattening the string values of the fields (“Department”, “Artist Display Name”, “Object Begin Date”, “Medium”, “Classification”, “Tags”) for the training set in the format as depicted in Listing 2, and second, in calculating, from the count of those values (per exhibition), the probability of each value showing in each exhibition, as depicted in the excerpt in the Listing 3.
"exhibitions": [
"x": "Title of exhibition is: Sculpture and Decorative Arts of the Spanish Renaissance and the description is: The Metropolitan Museum of Arts small but excellent collection of Spanish polychrome sculpture, including sacred reliefs and freestanding carved figures once housed in the churches of Spain, is displayed in the gallery adjacent to the newly reopened Velez Blanco Patio. The selection, which displays the unique blending of early western European and Islamic stylistic and technical influences, emphasizes the diversity in the material culture of Renaissance Spain after the Catholic reconquest by Ferdinand and Isabella.",
"y": {
"Department": [
"European Sculpture and Decorative Arts",
"European Sculpture and Decorative Arts",
"European Sculpture and Decorative Arts",
"The American Wing",
"European Sculpture and Decorative Arts",
"European Sculpture and Decorative Arts",
"European Sculpture and Decorative Arts",
"European Sculpture and Decorative Arts",
"European Sculpture and Decorative Arts",
"European Sculpture and Decorative Arts",
"European Sculpture and Decorative Arts"
"Artist Display Name": [
Listing 2: Flat version of the dataset for the Met Museum’s exhibitions, where $x$ is the title/description of the exhibition and $y$ is the flattened version of the metadata associated with the exhibition. We have therefore eliminated the artworks object ids.
"exhibitions": [
"x": "Title of exhibition is: Sculpture and Decorative Arts of the Spanish Renaissance and the description is: The Metropolitan Museum of Arts small but excellent collection of Spanish polychrome sculpture, including sacred reliefs and freestanding carved figures once housed in the churches of Spain, is displayed in the gallery adjacent to the newly reopened Velez Blanco Patio. The selection, which displays the unique blending of early western European and Islamic stylistic and technical influences, emphasizes the diversity in the material culture of Renaissance Spain after the Catholic reconquest by Ferdinand and Isabella.",
"y": {
"European Sculpture and Decorative Arts": 0.90909091,
"The American Wing": 0.09090909,
"Diego de Pesquera": 0.25,
"Juan Martinez Montanes": 0.25,
"Juan de Ancheta": 0.25,
"Diego de Atienza": 0.25,
"1600": 0.18181818,
"1500": 0.09090909,
"1585": 0.27272727,
"1630": 0.09090909,
"1567": 0.09090909,
"1615": 0.09090909,
"1575": 0.09090909,
"1646": 0.09090909,
"Tin-glazed earthenware": 0.09090909,
"Tin-glazed and luster-painted earthenware": 0.27272727,
"Silver gilt, enamel": 0.09090909,
"Wood, painted and gilt": 0.09090909,
"Wool, silk, metal thread on canvas": 0.18181818,
"Polychromed wood with gilding": 0.09090909,
"Wood, polychromed and gilded": 0.09090909,
"Silver gilt with enamel, cast, chased, and engraved": 0.09090909,
"Ceramics-Faience": 0.1,
"Ceramics-Pottery": 0.3,
"Sculpture": 0.3,
"Textiles-Embroidered": 0.2,
"Metalwork-Silver": 0.1
"z": [
"187702",
"187863",
"196434",
"197089",
"199674",
"210828",
"210826",
"201910",
"202718",
"205084",
"197090"
Listing 3: Probability flattening of the dataset, where now the $y$ variable represents the percentage of each metadata string occurring within the metadata for all the artworks associated with each particular exhibition, ignoring the metadata type – that is, we merge all metadata together for each exhibition. This allows a clear mapping between an exhibition title/description and a fixed-size high dimensional probability vector.
[…]
Figure 2: Neural network approach using self-contained text vectorisation and the (fixed-length) probability of finding a value on the exhibitions metadata fields. From left to right, the neural network takes a variable length plain text, i.e., the title and description of the exhibition, then it creates a vectorisation, which outputs a numeric vector, an embedding of that text. That vector has 256 dimensions, or numbers. Then it passes that vector to a one-dimensional average pooling layer, which averages those 256-dimensional vectors to smaller 64-dimensional vectors. That is passed to a so-called hidden layer, a dense layer of 256 nodes that has an activation function (in this case a ReLU function) – the ReLU (Rectified Linear Unit) function is defined as $\text{ReLU}(x)=\max(0,x)$, and outputs simply the input if it is positive, and zero otherwise. Finally, a second hidden layer with 8615 nodes or neurons and a linear activation function (in this case identity function) maps the results to the probabilities. In the diagram, $y_{\text{dim}}$ is the number of distinct values of the exhibitions’ artworks metadata fields (e.g., [“European Sculpture and Decorative Arts”, “The American Wing”, “Diego de Pesquera”, “1585”, “Sculpture”, …]). Therefore, $y_{\text{dim}}=8615$. The final outputs are probabilities of each one of those metadata fields (or what we call generalised tags) to show up in the list of artworks associated with that input title and description. We tested several activation layer functions, which decide the type of output. We also played with the number of hidden layer nodes, and settled on 256 nodes. The network/training parameters were: number of epochs = 2048, with the text vectorisation using max_tokens = 32768, output_sequence_length = 256, output_mode = "int", and standardize = "lower_and_strip_punctuation". For all runs we used batch_size = 16.
We run and test the model using several combinations made of different parameters and different modes of operation. The neural network we use is depicted, in schematic and compact fashion, in Figure 2. The data processing starts left to right: we feed in a title and description, the “xx” in Listing 3, and that text is cleaned and broken up; transformed into a vector; then after several layers, it outputs probabilities of each generalised tag to show up, and we shall use those later to map the most probable artworks from our database.
As described in the previous section, we divide our dataset in Table 1 into a training and a validation sets ($80\%$ examples for training (a total of 188 examples) and $20\%$ samples for validation or in-sample testing – a total of 48 examples, taking the total to the full set of 236 exhibitions. We train the model as described in the previous section, and monitor the error difference between the predicted probabilities of each generalised tag and the actual probabilities (within the training and validation sets). We display below in Figure 3 the loss or mean square error (MSE) for the training of the neural network for the entire training and validation sets.
[…]
Figure 3: Mean squared error (MSE) plotted for the training and validation sets of the neural network in Figure 2. The inset shows the same plot with a logarithmic y-axis.
From analysing Figure 3 we can conclude two key insights. First, that the models seem to converge as the MSE value decreases with epoch, as training progresses. Second, that it also seems to show that after 150 epochs or so, we start to have a serious case of over-fitting, as we can see the loss on the validation set stops decreasing, so the model is simply memorising the data after that.
We then attempted to analyse which of the generalised tag sub-groups is our model matching better (or worse). That is, for each exhibition, we look at the actual set of artworks on that exhibition and the predicted one, extract for each the generalised tags per sub-group (“Department”, “Artist Display Name”, “Object Begin Date”, “Medium”, “Classification”, “Tags”) and do a comparison, calculating the percentage of matching tags. We then average the percentage intersection or percentage matching across all exhibitions within the validation set (as within the training set this would not make sense, as we can have over-fitting, so we would not learn much by looking at those percentages.). So, e.g., if all the predicted artworks have as Department = “European Paintings” and the actual artworks the same, that percentage would be 1, a perfect match. Those results are depicted in Figure 4.
Figure 4: Percentage of generalised tags intersection on the validation set, between actual and prediction by the model.
The results show that the model is best at predicting the “Department” and “Classification” generalised tags, e.g., Department = “Asian Art”, “Photographs”, “Drawings and Prints”, “European Sculpture and Decorative Arts”, “Modern and Contemporary Art”, etc., and Classification = “Prints”, “Photographs”, “Paintings”, “Drawings”, “Creche”, etc. The model, however, is quite bad at getting the “Artist Display Name”, which is not surprising but it still disappointing. So, the model seems to focus more on the style and historical periods or traditions of the art exhibition. It is also quite disappointing that the model cannot achieve more than a $50\%$ matching percentage in any of the generalised tags at all.
The output is produced as follows. First, we run the model, obtaining an 8615-size numeric vector, like $y_{\text{output}}=[0.00566933,0.01132729,0.00091457,\ldots]$. These are the probabilities of each one of the possible values for the generalised tags within the full dataset of 236 exhibitions we have, e.g., […, ’Cups’, ’Curtains’, ’Cut Paper’, ’Cut and pasted fabrics and papers on paper’, …]. Both the output probability vector $p_{i}$ and the list of all generalised tags $\text{gtags}_{i}$ have a fixed size of 8615 items. We then calculate a hit score for each one of the rows of the full Met Museum objects or artworks database, which has a total of $484,956$ rows, with columns to be used being “Department”, “Artist Display Name”, “Object Begin Date”, “Medium”, “Classification” and “Tags”. These columns may have or not one of the generalised tags values $\text{gtags}_{i}$ above. The hit score $\text{hit}_{j}$ is calculated for row $j$ in the Met Museum database as:
$\text{hit}_{j}=\sum_{i=1}^{8615}p_{i}\cdot\delta_{ij}$ (1)
where
$\delta_{ij}=\begin{cases}1,&\text{if }\text{gtags}_{i}\text{ exists in row }j\\ 0,&\text{otherwise}\end{cases}$ (2)
We then sort the database in descending order of $\text{hit}_{j}$, and extract the top 16 artworks that represent the best matches. Given this, we can calculate the percentage of matched artworks, as depicted in Figure 5, is less than $1\%$. Even taken into account that we are trying to match a few artworks across a 400,000400,000+ artworks database, this is still very small. Taken together, the over-fitting, the low matching percentage of the generalised tags and the low matching of artworks seem to indicate the model is weak. We shall show that it is indeed very weak, next, when we test the model out-of-sample.
Figure 5: Percentage intersection of predicted artworks that match actual artworks on the validation set.
We tested the model out-of-sample and input the following title and description for a hypothetical exhibition. For this case, we used all the 236 exhibitions as training set, since we are doing out-of-sample prediction, there is no need to restrict ourselves by splitting the full dataset into a training and validation set, it would be wasteful. Notice also that the choice of the description was done on purpose since the word “impressionist” only shows up twice in a single exhibition, the Met Museum’s “Neo-Impressionism” [^11] This exhibition website page can be seen in https://www.metmuseum.org/exhibitions/listings/2001/paul-signac-circle. exhibition of 2001, with 7 paintings, and the sentence “still life” shows up 50 times within the generalised tags, well below the count for other generalised tags as we can see from Table 2. So, we are trying to make it difficult for the machine learning model to predict what generalised tags, and therefore what artworks to choose for our test input exhibition.
Title of exhibition is: The First Impressionists in Paris and the description is: the paintings of still life from the first Impressionists who created the modern art.
The results, using the full training set of 236 exhibitions with a training of 2048 epochs were as follows (results were as bad at 32 and 64 and 150 epochs too, before it starts to do over-fitting):
Object ID Department Artist Display Name Object Begin Date Medium Classification
38056 Asian Art [] 1630 Ink, opaque watercolor, and gold on paper [Paintings]
64533 Asian Art [] 1760 Ink, opaque watercolor, and gold on paper [Paintings]
56802 Asian Art [Torii Kiyonaga] 1742 Woodblock print; ink and color on paper [Prints]
67568 Asian Art [] 1800 Silk [Textiles-Woven]
37883 Asian Art [] 1620 Ink and opaque watercolor on paper [Paintings]
71999 Asian Art [] 1700 Silk / Compound weave [Textiles-Woven]
37812 Asian Art [] 1510 Ink and opaque watercolor on paper [Paintings]
37897 Asian Art [] 1640 Ink and opaque watercolor on paper [Paintings]
37920 Asian Art [] 1734 Ink and opaque watercolor on paper [Paintings]
50939 Asian Art [] 1600 Ink, opaque watercolor, and gold on paper [Paintings]
51680 Asian Art [Unidentified artist] 1800 Album leaf; ink and color on silk [Paintings]
671016 Asian Art [] 1500 Hanging scroll; ink, color and gold on silk [Paintings]
37174 Asian Art [Torii Kiyonaga] 1778 Woodblock print; ink and color on paper [Prints]
45034 Asian Art [Katsushika Hokusai] 1760 Woodblock print; ink and color on paper [Prints]
55258 Asian Art [Utagawa (Gountei) Sadahide] 1866 Triptych of woodblock prints; ink and color on paper [Prints]
59202 Asian Art [] 1800 Wood [Netsuke]
Figure 6: Marble composite of images related to the artworks from the Met Museum corresponding to the run with epochs=2048 for the model training with the entire dataset, using as input self-contained embeddings and output per exhibition metadata occurrence frequencies. Notice we only display the images of the artworks that are on the public domain, as provided by the Met Museum website and its API [The Metropolitan Museum of Art, 2025a].
The results, the artwork choices, which can be seen in the output above and in the marble composite image in Figure 6 show that the results are very poor. All the works are from “Asian Art” rather than “European Paintings”, and the date span is totally wrong, 1500–1866. This led us to attempt to either improve or choose another machine learning model for our task. Nonetheless, we decided to keep these results for comparison and to have a base (and basic) model. Our first improvement is the subject of the next section.
3.2 OpenAI-based embedding vectorisation
In this second approach, we use some external help in the form of the state-of-the-art OpenAI embedding model text-embedding-3-large that can perform very high-quality text vectorisation [OpenAI, 2024]. So, instead of relying exclusively on the text input (a total of $26,388$ words) and its statistics to create numeric vectors that hopefully will show some syntactic and semantic similarity and statistical analysis that can be used, we shall rely on the much larger training corpus of the OpenAI models, which are known to produce numeric vectors that encode a high-level of syntactic and semantic representation (see e.g., Keraghel et al. [2024], Giglietto [2024], Petukhova et al. [2025]). We apply this to the input text, the exhibition’s title and description (overview text) concatenation, replace the encoding that was done fully internally before, while keeping the rest of the neural network unchanged, as described in the previous methodology as depicted in Figure 2. The resulting network can be seen in Figure 7 below.
[…]
Figure 7: Neural network approach using OpenAI embedding model text-embedding-3-large and the (fixed-length) probability of finding a value on the exhibitions metadata fields.
One of the advantages of using the OpenAI embedding models to create numeric vectors from the input text is that these models use the so-called Transformer architecture [Vaswani et al., 2017]. These models take into account not only the word or token frequency, but also the direction or order of the text (so “…will many…” is encoded differently from “…many will…”) and contextual information such as recognising and recording that the “he” in “…the direction of Picasso’s work changed, and he moved away from Cubist approaches…” is related to “Picasso”. These models have semantic encoding, recording not just the probability distribution function of words but also the meaning, context and direction and far-away relationships. Furthermore, these models have been trained in a vast corpus of textual data, and therefore have better vector similarity, among other advantages, for semantically similar word sequences, particularly in English language.
[…]
Figure 8: Mean squared error (MSE) plotted for the training and validation sets for the neural network in Figure 7, with an inset zoomed region.
As before, the first thing to do is to train the model, we use the same training and validation sets as before ($80\%$ of examples for training – a total of 188 examples, and $20\%$ samples for validation or in-sample testing – a total of 48 examples). We optimise the model using the same numerical approach as before (Adam optimisation). We measure the same mean squared error (MSE) as before, i.e., the sum of squares of the differences in predicted versus actual probabilities of occurrence of each one of the generalised tags (we recall that there are $8591$ unique number of those). The MSE values for the training and validation sets as a function of the number of training epochs is depicted in Figure 8.
The results seem to show a slightly better convergence on both the training and validation sets, when one compares with the MSE of the previous model in Figure 3. The MSE in Figure 8 seems to show we reach convergence around 50 training epochs or so. We can also see, as depicted in Figures 9 and 10, that as the training progresses, the percentage of tags that it gets right increases. Here the results seem much of an improvement with respect to the previous model, if we compare against the results in Figure 4.
Figure 9: Percentage of generalised tags intersection on the validation set, between actual and prediction by the model.
Figure 10: Percentage of generalised tags intersection on the validation set, between actual and prediction by the model.
The results show that, at 64 epochs, the percentage of intersection for each of the generalised tag sub-groups is higher for this model with OpenAI embedding than for the one with self-contained embeddings (previous section). Furthermore, we can see the results are stable for longer runs, as can be seen in the run at 2048 epochs, in Figure 10. In fact, for “Department” and “Classification”, the matching is around $70\%-80\%$, which is encouraging. We also analyse the average percentage of intersection of predicted artworks per exhibition versus actual artworks within the validation set, and this is depicted in Figure 11.
Figure 11: Average percentage of predicted artworks that match actual artworks on the validation set. Note that the percentage overlap based on random choice is around $0.00907\%$.
The results seem to show also an improvement with respect to the first model, the base model with self-contained embeddings (depicted in Figure 5). The results jump from around $0.35\%$ to around $0.8\%$. In fact, when we increase the number of epochs to 2048, we get a convergence and saturation of this intersection value at around $1.34\%$. Although this value seems small, one has to think carefully how much larger this value is with respect to the value one obtains if we select randomly $k$ artworks from the full N=484,956N=\$484,956\$ works of art dataset. If we draw randomly $k=44$ ($44.36$ is the average number of artworks across all our 236 exhibitions) artworks from the full dataset, the average count of artworks expected overlap or intersection is $k\times\frac{k}{N}=\frac{k^{2}}{N}=\frac{44^{2}}{484956}=\approx 0.00399$. So, the percentage overlap (count of overlap divided by the number of artworks in each exhibition) can be approximated by $k\times\frac{k}{k\times N}=\frac{k}{N}=\frac{44}{484956}=\approx 9.072988\times 10^{-5}\approx 0.00907\%$. So, for our current model we can get a accuracy 150 times or so better than random choice (even for our basic self-contained model we get an accuracy 70 better than random choice). The question now is, do these better results reflect themselves on better out-of-sample predictions, where these models should really count?
We test this by again running our model out-of-sample, using as training set all 236 example exhibitions we have, and inputting the following title and description for a hypothetical exhibition.
Title of exhibition is: The First Impressionists in Paris and the description is: the paintings of still life from the first Impressionists who created the modern art.
As we have done before, we run the model to obtain a fixed-size output probability vector $y_{\text{output}}$, again of length 8615, where each element $p_{i}$ corresponds to the probability of each of the generalised tags in $\text{gtags}_{i}$. This is for all generalised tags observed across all 236 exhibitions. We then compute the hit score as before, given by the formula (1). Finally, we sort all entries by $\text{hit}_{j}$ in descending order and select the top 16 artworks with the highest hit scores. The output was, using this time 128 epochs for training:
Object ID Department Artist Display Name Object Begin Date Medium Classification
436526 European Paintings [Vincent van Gogh] 1890 Oil on canvas [Paintings]
436525 European Paintings [Vincent van Gogh] 1890 Oil on canvas [Paintings]
436528 European Paintings [Vincent van Gogh] 1890 Oil on canvas [Paintings]
437480 European Paintings [Hubert Robert] 1777 Oil on canvas [Paintings]
437984 European Paintings [Vincent van Gogh] 1889 Oil on canvas [Paintings]
437654 European Paintings [Georges Seurat] 1887 Oil on canvas [Paintings]
437658 European Paintings [Georges Seurat] 1884 Oil on canvas [Paintings]
436534 European Paintings [Vincent van Gogh] 1890 Oil on canvas [Paintings]
435849 European Paintings [Carolus-Duran (Charles-Auguste-Emile Durant)] 1890 Oil on canvas [Paintings]
435868 European Paintings [Paul Cézanne] 1890 Oil on canvas [Paintings]
436923 European Paintings [Maximilien Luce] 1890 Oil on canvas [Paintings]
436529 European Paintings [Vincent van Gogh] 1888 Oil on canvas [Paintings]
436531 European Paintings [Vincent van Gogh] 1885 Oil on canvas [Paintings]
436536 European Paintings [Vincent van Gogh] 1889 Oil on canvas [Paintings]
438722 European Paintings [Vincent van Gogh] 1885 Oil on canvas [Paintings]
437378 European Paintings [Odilon Redon] 1890 Oil on canvas [Paintings]
Figure 12: Marble composite of images from the Met Museum corresponding to the run with epochs=128 for the model training with the entire dataset, using as input OpenAI embeddings and output per exhibition metadata occurrence frequencies. Notice we only display the images of the artworks that are on the public domain, as provided by the Met Museum website and its API [The Metropolitan Museum of Art, 2025a]. Non-available images are marked by their object id JPEG name.
Also, the output was, using the longer 2048 epochs training:
Object ID Department Artist Display Name Object Begin Date Medium Classification
436525 European Paintings [Vincent van Gogh] 1890 Oil on canvas [Paintings]
436528 European Paintings [Vincent van Gogh] 1890 Oil on canvas [Paintings]
437378 European Paintings [Odilon Redon] 1890 Oil on canvas [Paintings]
436526 European Paintings [Vincent van Gogh] 1890 Oil on canvas [Paintings]
436534 European Paintings [Vincent van Gogh] 1890 Oil on canvas [Paintings]
436530 European Paintings [Vincent van Gogh] 1888 Oil on canvas [Paintings]
437984 European Paintings [Vincent van Gogh] 1889 Oil on canvas [Paintings]
435849 European Paintings [Carolus-Duran (Charles-Auguste-Emile Durant)] 1890 Oil on canvas [Paintings]
435868 European Paintings [Paul Cézanne] 1890 Oil on canvas [Paintings]
435880 European Paintings [Paul Cézanne] 1890 Oil on canvas [Paintings]
436923 European Paintings [Maximilien Luce] 1890 Oil on canvas [Paintings]
847023 European Paintings [Frederic, Lord Leighton] 1890 Oil on canvas [Paintings]
437654 European Paintings [Georges Seurat] 1887 Oil on canvas [Paintings]
437655 European Paintings [Georges Seurat] 1881 Oil on canvas [Paintings]
437658 European Paintings [Georges Seurat] 1884 Oil on canvas [Paintings]
438015 European Paintings [Georges Seurat] 1886 Oil on canvas [Paintings]
These results, in particular the second one, with a longer training (even if it over-fits in the training set), are more in line with our expectations that our model is now working better as intended. We can also see that from the marble composite depicted in Figure 13. The first element that seems to improve was that we now have a much tighter artwork creation interval, all artworks are within the period 1777–1890, a span of 113 years for the run with 128 epochs and 1881–1890, a span of only 9 years, for the longer run with 2048 epochs. Impressionism is considered to be a period of art history around 1860–1886 [Herbert, 1988, Gombrich, 2023], although this is not exactly fixed, obviously. Nonetheless, our results are close, even if our system seems to have an inclination to choose a lot of Post-Impressionists and Neo-Impressionists rather than actual from the first Impressionists, as we asked it to do by using as input “…First Impressionists in Paris…”. The second element that seems to also improve, is that now we have all artworks “Department” = “European Paintings” as opposed to “Asian Art”, as it should be, and “Classification” as “Paintings”, not as “Prints” or “Textiles-Woven”. The model also seems to make an effort to find instances of still life, as can be seen by the 4 matches in Figure 12 and the 5 matches in Figure 13 – not only that is a good percentage of matches, but increases with increasing number of training epochs, as it should. [^12] We note that we can examine freely all the images of all the artworks, even if we cannot display them in this article, by going to the URL https://www.metmuseum.org/art/collection/search/xxxxxx, where xxxxxx is the object_id. This way we can decide, by visual inspection, which one are instances of still life or not.
Figure 13: Marble composite of images from the Met Museum corresponding to the run with epochs=2048 for the model training with the entire dataset, using as input OpenAI embeddings and output per exhibition metadata occurrence frequencies. Notice we only display the images of the artworks that are on the public domain, as provided by the Met Museum website and its API [The Metropolitan Museum of Art, 2025a]. Non-available images are marked by their object id JPEG name.
The results are also interesting in the sense that only $5$ out of the predicted $16$ artworks, $31.25\%$ of them, are present in our full dataset of 236 exhibitions, meaning that the model seems to be able to select artworks outside the training set, and select from the full dataset of 400,000+\$400,000\$+ artworks in the Met Museum. This was one of the objectives of this work, to be able to generalise outside the given input exhibition dataset. Therefore, we seem to be going in the right direction by using more advanced embeddings. In the next section, we go one step further in this direction.
3.3 Direct neural network embedding output to embedding metadata of artworks
Given that the OpenAI embeddings improved the accuracy of our model, we decided to use it not just for the input but for both the inputs and outputs, i.e., the text/descriptions of the exhibitions and the corresponding concatenated generalised tags for all the artworks in each exhibition. [^13] We note that there are several ways to encode the output target values, the generalised tags. One alternative would be to encode each exhibition against each of the generalised tags present in each of the subgroups, or one-to-one even, in effect creating and unfolding more training examples. We decided not to complicate this model too much and simply try to map the encoding or embedding of the title description against the encoding or embedding or a single long string corresponding to all the concatenated generalised tags for all the artworks in each exhibition in the order they show up on the exhibition webpage and the Met Museum database. We also create the embeddings for all the Met Museum’s $484,956$ works of art, for their corresponding generalised tags or metadata and store those on a vector database [Andoni and Indyk, 2008]. The reasons we use a vector database are: first, the embeddings data structure becomes too large for our computational resources/capacity available; second, it becomes too slow to search for the best exact nearest neighbours in an array in memory/disk; and third, that these vector databases are optimal for storing embeddings and provide already built-in mechanisms for nearest neighbour search [Iscen et al., 2017], which will be the approach we take in order to find the artworks (or metadata embeddings) corresponding to the input texts (or title/description embeddings). We then train the model and when that trained model is passed an input text (either from the training or validation sets or from a new out-of-sample test set), it creates an output embedding vector, for which we can then find the nearest neighbours on the vector database of all Met Museum’s artworks embeddings. Those will be the selected artworks for that corresponding exhibition. We use the Facebook AI Similarity Search (FAISS) library as described in Douze et al. [2024], in particular we use the Inverted File with Flat (IndexIVFFlat) vectors model, since the exact match approach could not fit the memory available on our modest computational resources. Below in Figure 14 we can find a schematic of the neural network for this kind of modelling.
[…]
Figure 14: Neural network approach using the OpenAI embedding model text-embedding-3-large for both the inputs: the title/descriptions of the exhibitions and the outputs, the concatenated flattened list of generalised tags for the exhibitions.
The main difference to the previous model is that now we use the two embeddings, in two directions. Once we train the neural network to map the input embeddings to output embeddings, we can then predict output embedding(s) for training/validation/test inputs. We then use those output embedding(s) to search for $k$ nearest embeddings neighbours (where $k$ is the number of artworks on the input exhibition, or $k=16$ if not available). Those $k$ nearest embeddings neighbours correspond to $k$ nearest artworks from our large Met Museum database. The training follows the same approach as before, with the training and validation sets split in the same way. The MSE for the model was as follows, as depicted in Figure 15.
Figure 15: Mean squared error (MSE) plotted for the training and validation sets for the neural network in Figure 14.
The convergence is again quite good, with the saturation for the validation set around 120–150 epochs. This is also reflected in the chart of the intersection percentage of the generalised tags per sub-group as a function of epoch within the validation set in Figure 16.
Figure 16: Intersection percentage of the generalised tags per sub-group as a function of epoch within the validation set for the neural network in Figure 14.
The results for the percentage of intersection predicted versus actual artworks for the validation set is depicted below in Figure 17. It is higher (at 2048 epochs) than for the models before. We get, as can be seen Figure 17, of around $4\%$ or below. The value at 2048 epochs is $3.72\%$, so around 400 times better than random choice.
Figure 17: Average percentage of intersection artworks for the validation set for the neural network in Figure 14. Again, note that the percentage overlap based on random choice is around $0.00907\%$, as demonstrated in the text.
We then test the model out-of-sample, and input the following title and description for a hypothetical exhibition.
Title of exhibition is: The First Impressionists in Paris and the description is: the paintings of still life from the first Impressionists who created the modern art.
The output was, using 2048 epochs and a number of probes $n_{probe}=4$ (notice that probes refers to the number of clusters that the algorithm used by FAISS search use during a query – the higher the more accuracy we get but less performance/speed):
Object ID Department Artist Display Name Object Begin Date Medium Classification
693862 Drawings and Prints [Goupil et Cie, Braun & Cie, …] 1914 Photogravure [Prints]
356848 Drawings and Prints [Edward Ancourt, Gustave Courbet, …] 1868 Lithograph with hand coloring [Prints]
489100 Modern and Contemporary Art [Jean Metzinger] 1917 Oil on canvas [Paintings]
437303 European Paintings [Camille Pissarro] 1880 Oil on canvas [Paintings]
488853 Modern and Contemporary Art [Edouard Vuillard] 1889 Oil on canvas [Paintings]
679686 European Paintings [Fernand Khnopff] 1884 Oil on canvas [Paintings]
483504 Modern and Contemporary Art [Georges Braque] 1923 Oil and sand on canvas [Paintings]
436529 European Paintings [Vincent van Gogh] 1888 Oil on canvas [Paintings]
435997 European Paintings [Pierre-Auguste Cot] 1880 Oil on canvas [Paintings]
388319 Drawings and Prints [Jules-Edmond-Charles Lachaise, …] 1850 Oil paint on canvas support [Drawings]
438435 European Paintings [Claude Monet] 1872 Oil on canvas [Paintings]
437347 European Paintings [Pierre Puvis de Chavannes] 1878 Oil on canvas [Paintings]
490035 Modern and Contemporary Art [Edouard Vuillard] 1897 Oil on cardboard [Paintings]
438136 European Paintings [Paul Cézanne] 1870 Oil on canvas [Paintings]
435876 European Paintings [Paul Cézanne] 1888 Oil on canvas [Paintings]
482211 Modern and Contemporary Art [Jacques Villon] 1905 Oil on canvas [Paintings]
The marble image for these works of art is depicted below in Figure 18. The results in the above table and displayed in Figure 18 seem to show an improvement. In particular, we now find some early impressionists, such as Claude Monet and Camille Pissarro. We also have a high percentage of intersection of predicted artworks versus actual artworks on the validation set, as depicted in Figure 17, with an end value at 2048 epochs of around 400 times larger than the one expected by random choice. In terms of finding instances of still life themes, we find that we get only two still life instances, and furthermore these are from a cubist, Jean Metzinger, a painting from 1917, and from a symbolist, Fernand Khnopff, with a painting from 1884, as can be seen in Figure 18. Finally, the time interval of the predicted artworks, 1850–1923, 73 years, is much larger than the length of the impressionist period. So, this does not seem like a conclusive improvement over the previous simpler model, the one in Section 3.2. We now attempt using the full-scale power of OpenAI LLM models in the next section.
Figure 18: Marble composite of images from the Met Museum corresponding to the run with epochs = 2048 for the model training with the entire dataset, using as input OpenAI embeddings of the title/description of each exhibition and output the OpenAI embeddings of concatenated generalised tags for all artworks in each exhibition. Notice we only display the images of the artworks that are on the public domain, as provided by the Met Museum website and its API [The Metropolitan Museum of Art, 2025a]. Non-available images are marked by their object id JPEG name.
3.4 Full OpenAI approach including embedding and mapping
Fourth and finally, we fully use external help, this time by using one of the latest OpenAI GPT (Chat Generative Pre-Trained Transformer models – Radford et al. [2018], Brown et al. [2020]), in this case gpt-4o-mini. These GPT-based models are trained from very large amounts of data scraped from the open Internet, in particular Wikipedia, the arXiv.org e-Print archive, Reddit, among others. These models are a subset of the so-called large language models (LLMs), whereby one maps an input variable $x$, corresponding to the question, to an output variable $y$, corresponding to the answer, using a complex mathematical function, which is encoded in the values of the weights of the nodes of a (very large [^14] It is estimated that this particular model may have up to 8 billion parameters. However, this is just an estimate as OpenAI usually does not officially disclose exact number of parameters in their models.) neural network. The transformation of the question, which is a plain text input, into a numerical vector $x$ of very high dimension is similar to the method described in the Sections 3.1, 3.2 and 3.3, an embedding, but it is assumed (since again it is not disclosed by OpenAI) that gpt-4o-mini has its own embedding model. The same goes for the output variable $y$. The mapping is not an explicit analytical equation, but rather a numerical function discovered by the training process using a technique called back-propagation [Rumelhart et al., 1986], and encoded in the numerical weights of the nodes of a very large neural network. In order to use gpt-4o-mini to create a list of artworks for an exhibition, we use the following steps. First, we fine-tune the model, by showing it examples in the form $\langle x,y\rangle$, where $x$ is the input text as in the listing above 2 and $y$ is a table of the metadata items for all the artworks presented within the exhibition (see Listing 4 for an example). Fine-tuning is a machine learning technique whereby one takes an already pre-trained model (such as one of the GPT models) and adapts it to a new domain and/or makes it more specific to a sub-domain [Yosinski et al., 2014, Howard and Ruder, 2018, Radford et al., 2018, Wei et al., 2021]. The fine-tuning process re-trains the last few layers of the neural network, while maintaining most layers intact. This process then improves the model for the new domain and/or sub-domain we are interested in. [^15] It is well known that there is a kind of feature hierarchy in neural networks, whereby the first layers encode small features while the last layers encode overall generic task or domain specific features, as demonstrated in Yosinski et al. [2014]. A common analogy used is of a very young child learning a second language, the knowledge gained during the first language acquisition is re-used and fine-tuned for the second language. An example would be to retrain the weights between the second hidden layer and the output layer on our example toy model neural network in Figure 1.
{"messages":
{"role": "system", "content": "ArtCurator is a factual chatbot that is an expert in JSON format and in artworks and exhibitions from The Metropolitan Museum of Art."},
{"role": "user", "content": "Title of exhibition is: Sculpture and Decorative Arts of the Spanish Renaissance and the description is: The Metropolitan Museum of Arts small but excellent collection of Spanish polychrome sculpture, including sacred reliefs and freestanding carved figures once housed in the churches of Spain, is displayed in the gallery adjacent to the newly reopened Velez Blanco Patio. The selection, which displays the unique blending of early western European and Islamic stylistic and technical influences, emphasizes the diversity in the material culture of Renaissance Spain after the Catholic reconquest by Ferdinand and Isabella."},
{"role": "assistant", "content": "\"{’Department’: [’European Sculpture and Decorative Arts’, European Sculpture and Decorative Arts’, European Sculpture and Decorative Arts’, The American Wing’, European Sculpture and Decorative Arts’, European Sculpture and Decorative Arts’, European Sculpture and Decorative Arts’, European Sculpture and Decorative Arts’, European Sculpture and Decorative Arts’, European Sculpture and Decorative Arts’, European Sculpture and Decorative Arts’], Artist Display Name’: [’None’, None’, None’, None’, Diego de Pesquera’, None’, None’, None’, Juan Martinez Montanes’, Juan de Ancheta’, Diego de Atienza’], Object Begin Date’: [’1600’, ’1500’, ’1585’, ’1630’, ’1567’, ’1585’, ’1585’, ’1600’, ’1615’, ’1575’, ’1646’], Medium’: [’Tin-glazed earthenware’, Tin-glazed and luster-painted earthenware’, Tin-glazed and luster-painted earthenware’, Silver gilt, enamel’, Wood, painted and gilt’, Wool, silk, metal thread on canvas’, Wool, silk, metal thread on canvas’, Tin-glazed and luster-painted earthenware’, Polychromed wood with gilding’, Wood, polychromed and gilded’, Silver gilt with enamel, cast, chased, and engraved’], Classification’: [’Ceramics-Faience’, Ceramics-Pottery’, Ceramics-Pottery’, None’, Sculpture’, Textiles-Embroidered’, Textiles-Embroidered’, Ceramics-Pottery’, Sculpture’, Sculpture’, Metalwork-Silver’]}\""}
Listing 4: The excerpt of the dataset transformed for the purpose of fine-tuning the GPT model, showing the same exhibition by the Met Museum, as in the previous listings, displaying the role of the GPT model we want it to adopt, the input or user content and the expected GPT model answer, the assistant content.
The model used can be pictured as in Figure 19 below. We fine-tune the model with the examples from the training set (we split the 236 exhibitions in the same way as in previous sections) in the format as in Listing 4. The fine-tuning includes checking the validation set, which is used only for monitoring and evaluation purposes (no automated hyper-parameter optimisation). The training set is used to create a fine-tune GPT model, which can then be called to do evaluation both in the training and validation sets, or on a test set made of out-of-sample examples.
Titles andoverview texts+artworks and metadataFine-tuningOpenAIgpt-4o-miniFine-tunedmodelTitle andoverview textFine-tunedmodelPredictedArtworks andmetadata
Figure 19: Two-step neural network approach using the fine-tune versions of OpenAI’s gpt-4o-mini (version 2024-07-18). The input text is the title/description of each exhibition, and the output is a formatted version of the listing 4.
After we fine-tune the model, in this case gpt-4o-mini (version 2024-07-18), we can use it directly. We can query using inputs within the validation set or out-of-sample test set examples. As before, we split the set of exhibitions we have into a training and a validation set, in the proportion $80\%$ examples for training (a total of 188 examples) and $20\%$ samples for validation or in-sample testing (a total of 48 examples). We notice that the output of the model this time can be of variable length, that is to say, the number of predicted artworks this time is variable (in all the other models before, the selection is done based on the number of actual artworks in each exhibition). This we see as an advantage of this model. We also notice that more often than not the model outputs a series of artworks with generalised tags in an invalid format (e.g., values missing, rows/columns incomplete, etc.). In these, we simply rerun the model until we get a output that is fully capable of being parsed, having given up trying to fix the invalid formats. We also notice that the output of GPT although it is a list of rows with “Department”, “Artist Display Name”, “Object Begin Date”, “Medium”, “Classification”, “Tags”, these are not necessary either real existing artworks (i.e., inconsistent data can be output, the so-called LLM hallucinations; see e.g., Farquhar et al. [2024] and references therein) or artworks within the Met Museum. Therefore, we still need to use the Met Museum database to find the nearest artworks to our output, using the same approach with the hit function as defined in Equation 1.
We plot the MSE of the training and validation sets for our fine-tuning in Figure 20. This time the convergence was not as clear as in the other models, however, this was not a problem as we can see later – fine-tuning typically changes the model weights only slightly rather than completely refitting the training dataset, so loss (MSE) reduction can be modest, see e.g., Howard and Ruder [2018, Fig. 4, p. 336].
Figure 20: Mean squared error (MSE) plotted for the training and validation sets for the neural network in Figure 19. The model was obtained from fine-tuning gpt-4o-mini-2024-07-18, with parameters batch_size=16 and learning_rate_multiplier=0.3 and the total number of trained tokens was $3,036,784$.
While before we had the charts of percentage of generalised tags intersection on the validation set, between actual and prediction by the model and the average of the percentage of predicted artworks that match actual artworks, we notice that OpenAI does not provide the intermediate fine-tuned models at each epoch (only at the end), so we could not run those metrics on all intermediate steps. The results at the end, using the validation set are:
Generalised tag
intersection (%)
Predicted artworks
matching actual (%)
Self-contained text
vectorisation
Department 50.35%
Artist Display Name 3.43%
Object Begin Date 14.85%
Medium 11.88%
Classification 48.21%
Tags 18.83%
0.637%
OpenAI-based embedding
vectorisation
Department 80.17%
Artist Display Name 3.28%
Object Begin Date 20.18%
Medium 24.68%
Classification 68.28%
Tags 26.86%
1.336%
Direct NN embedding
to artwork metadata
Department 77.85%
Artist Display Name 2.90%
Object Begin Date 19.92%
Medium 18.01%
Classification 42.98%
Tags 25.65%
3.720%
Full OpenAI approach
(embedding + mapping)
Department 88.19%
Artist Display Name 42.88%
Object Begin Date 31.89%
Medium 40.47%
Classification 76.22%
Tags 22.26%
7.623%
Table 5: Comparison of model performance on the validation set based on generalised tag intersection and actual artwork prediction accuracy. For the first, second and third models, we use epochs = 2048, while for the fourth model we use the fine-tuned model as described in the text in Section 3.4. Again, note that the percentage overlap based on random choice is around $0.00907\%$, as demonstrated in the text.
The results seem to show an improvement in all the generalised tag intersection percentages as well as the percentage of intersection of predicted versus actual artworks. The latter is around $7.623\%$, which is around 800 or so times what one could expect if we did a random choice.
As before, we also test the model out-of-sample, and input the following title and description for a hypothetical exhibition.
Title of exhibition is: The First Impressionists in Paris and the description is: the paintings of still life from the first Impressionists who created the modern art.
The output was, using the fine-tuned model:
Object ID Department Artist Display Name Object Begin Date Medium Classification
437299 European Paintings Camille Pissarro 1867 Oil on canvas Paintings
436945 European Paintings Edouard Manet 1862 Oil on canvas Paintings
437317 European Paintings Camille Pissarro 1872 Oil on canvas Paintings
436020 European Paintings Gustave Courbet 1873 Oil on canvas Paintings
436961 European Paintings Edouard Manet 1864 Oil on canvas Paintings
436952 European Paintings Edouard Manet 1867 Oil on canvas Paintings
436951 European Paintings Edouard Manet 1862 Oil on canvas Paintings
436964 European Paintings Edouard Manet 1866 Oil on canvas Paintings
436013 European Paintings Gustave Courbet 1873 Oil on canvas Paintings
437303 European Paintings Camille Pissarro 1880 Oil on canvas Paintings
436960 European Paintings Edouard Manet 1866 Oil on canvas Paintings
438002 European Paintings Edouard Manet 1880 Oil on canvas Paintings
438009 European Paintings Berthe Morisot 1865 Oil on canvas Paintings
436002 European Paintings Gustave Courbet 1866 Oil on canvas Paintings
436001 European Paintings Gustave Courbet 1865 Oil on canvas Paintings
435965 European Paintings Camille Corot 1865 Oil on canvas Paintings
437104 European Paintings Claude Monet 1864 Oil on canvas Paintings
436150 European Paintings Edgar Degas 1867 Oil on canvas Paintings
441111 European Paintings Michele Gordigiani 1864 Oil on canvas Paintings
437436 European Paintings Auguste Renoir 1865 Oil on canvas Paintings
441104 European Paintings Auguste Renoir 1880 Oil on canvas Paintings
437097 European Paintings Jean-François Millet 1869 Oil on canvas Paintings
437094 European Paintings Jean-François Millet 1872 Oil on canvas Paintings
437107 European Paintings Claude Monet 1876 Oil on canvas Paintings
441351 European Paintings Georges Clairin 1872 Oil on canvas Paintings
436519 European Paintings Pierre-Paul-Léon Glaize 1873 Oil on canvas Paintings
437111 European Paintings Claude Monet 1880 Oil on canvas Paintings
437109 European Paintings Claude Monet 1880 Oil on canvas Paintings
437110 European Paintings Claude Monet 1880 Oil on canvas Paintings
435627 European Paintings Charles-Edouard de Beaumont 1875 Oil on canvas Paintings
441353 European Paintings Benjamin-Constant (Jean-Joseph-Benjamin Constant) 1876 Oil on canvas Paintings
435979 European Paintings Camille Corot 1862 Oil on canvas Paintings
437133 European Paintings Claude Monet 1867 Oil on canvas Paintings
437135 European Paintings Claude Monet 1869 Oil on canvas Paintings
437138 European Paintings Claude Monet 1880 Oil on canvas Paintings
435984 European Paintings Camille Corot 1865 Oil on canvas Paintings
Figure 21: Marble composite of images from the Met Museum corresponding to the run for the fine-tune model trained on the training set only, using as input the text (title/description) of the exhibition and as output the actual list of artworks and their metadata in their original format as in the listing 4. Notice we only display the images of the artworks that are on the public domain, as provided by the Met Museum website and its API [The Metropolitan Museum of Art, 2025a]. Non-available images are marked by their object id JPEG name.
The results in the above table and in Figure 21, seem to show some, but not all, improvements across all metrics. First, the list contains more actual impressionists: Camille Pissarro, Claude Monet, Edgar Degas, Auguste Renoir and Berthe Morisot, so the best result so far. Second, we find 3 instances of still life:“Still Life with Apples and Pitcher” by Camille Pissarro, a painting from 1872; “Peonies” by Edouard Manet from 1864 and “Still Life with Flowers and Prickly Pears” by Auguste Renoir from around 1880 or later – so a result less impressive than the one seen in the model in Section 3.2, which had 5 instances of still life artworks, and on a smaller predicted set. Third, and finally, the predicted artworks encompass the period span of 1862–1880, 18 years in total, which is a good match to the actual Impressionist period.
4 Possible further research approaches
In this article we have focused on four specific approaches, in the above sections, to creating an AI-based curated exhibition. Nonetheless, there are many other ways this research can be expanded and/or enhanced. Here we suggest and review some of those possible avenues of research.
In terms of data, we can think of three main expanded approaches. The first approach one might consider is using more data, from the same data source we used, i.e., the Met Museum. As described in Table 1 we have managed to download a total of 1009 exhibitions from the Met Museum, but we have only used 236 of those. In order to increase this number, we suggest one could do the following:
  1. 1.
    Collect more input text (the exhibitions’ overview descriptions) by hand, as we have automated (via Python) the download of the relevant text. This meant some of them had problems (different formatted text, non-English text, non-ASCII text, image/table interference, etc.) and were removed automatically;
  2. 2.
    Some of the exhibitions were removed because our Python code could not find or could not parse the list of artworks associated with those exhibitions, in particular the artworks object ids which map to the Met Museum’s database. In principle one could recover some of those removed exhibitions by doing this work by hand;
  3. 3.
    The Python code was designed so that it could find the tabs/web page fields related to the title, overview text and exhibition objects, but we have noticed the Met Museum have changed several times their web design, not surprisingly, given our dataset spans over a period of 20+ years. Again, one could make an additional effort and parse those rejected or ignored exhibition pages by doing the work by hand.
Second, we could also enhance the number of usable exhibitions by adding those that are composed of artworks that are lent to the Met Museum. However, the reason those were removed is that the Met Museum database does not contain those artworks or theirs associated metadata. To manually add these would be a possible but very difficult and time-consuming task. Third, we could also gather data from other museums, and merge with the dataset we already have from the Met Museum. There are many possibilities, e.g., the Museum of Modern Art [The Museum of Modern Art, 2025a] provides a collection API; the British Museum has an extensive collection of more than five million objects [British Museum, 2025]; the National Gallery in London also provides an API [The National Gallery, 2025]; and many others. However, this would be a much larger task, and one would have to use as input not just the exhibition test text but also the target museum, as one would need constraints to minimise artworks hosted outside this target museum, to make the exhibition as feasible as possible. Furthermore, each museum uses its own format for the artwork metadata, with different field names/meanings, making the merging of data across multiple sources a very difficult and subjective task.
In terms of modelling itself, we can also think of four different approaches we could use. First, as mentioned in Section 3.1 we could use a Long Short-Term Memory (LSTM) to turn our task into a variable input/variable output problem. This is similar to what is used in human languages translations, e.g., English text to a French text [Sutskever et al., 2014]. The advantage of a LSTM neural network is that being a recurrent neural network it would allow for variable input/output sizes, which is part of our problem. While we have gone around this issue in this article by constructing models that transform the problem into a fixed input/output size, there is in principle no actual obstacle in using a LSTM to output a variable list of artworks. However, in practice with such a small amount of training examples (maximum of 236), the data size made this approach unrealistic. Maybe with a much larger dataset, including other museum’s data would make it possible. Second, we could construct a slightly different approach as we used in model in Section 3.3. While in that model we created an embedding of the input exhibition title/description and an embedding of the output concatenated list of metadata generalised tags (one string), we could instead unfold our training dataset by mapping each exhibition’s embedding to each metadata string embedding, so e.g., if one exhibition had $n$ artworks each with $m$ generalised tags, then we would create in effect $nm$ input/output embeddings. The reason we did not try this is that the resulting graph database and subsequent search would be too large/slow for our purpose, given our computational constrains. Third, we could in principle add our own tags by downloading all images of each artwork from the Met Museum database and using a Convolutional Neural Network (CNN) to extract our own extra features/metadata [LeCun et al., 2002, Krizhevsky et al., 2012]. However, this is a massive undertaking, requiring lots of computational power and storage space. Furthermore, this would require inside access as most images are not in the public domain. Of the total $484,956$ works of art, only $248,472$ are in the public domain and available for download. Fourth, and finally, we could also include in our models more of the Met Museum generalised tags. There are 50+ fields that the Met Museum provides in its full artwork metadata set, we just use 6 of those, as we recall these were “Department”,“Artist Display Name”, “Object Begin Date”, “Medium”, “Classification”, “Tags”. We did attempt to add other fields, but saw such minor or null improvements, at a computational cost and reduction of simplicity/understanding which we deemed unnecessary.
5 Conclusion
We have attempted to create four machine learning or artificial intelligence models that learn the behaviour of human art curators. We used data from 25 years of exhibitions at the Metropolitan Museum of Art in New York to train and optimise the models. Three of the four models can be shown to imitate human behaviour to a reasonable accuracy, namely, well above random choice for the set of existing exhibitions and the set of existing artworks in the full collection. Furthermore, as the models get increasingly complex, i.e., as we move from relying simply on the dataset statistical information and start using more and more the knowledge inherent in the so-called large language embedding and query models, we can show we get some improvement in the accuracy, measured by the selection of artworks and its metadata. Nonetheless, the improvement is not very large, and this seems to show that even models with modest size (e.g., our second model with input embeddings and output metadata probabilities) can perform almost as well as very large models using brute force. In particular, this can be seen in our out-of-sample experiment (the one with the impressionists related input text). Finally, we believe that as we gather more data in the future and improve the models, that these kinds of machine learning approaches may be able to get closer and closer to the level of curatorship that humans possess, although it remains a subject of discussion in philosophy of art if there is any artistic value in this. This could be of interest to art galleries who wish to use AI models to autonomously curate their in-house exhibitions and/or to allow the public to access a website or an app that would create virtual exhibitions based on the public’s input exhibition title and description.
Declarations
  • Funding: The author and the research presented in this article were fully and wholly self-funded by the author;
  • The author declares no existing conflicts of interest;
  • Availability of the dataset and materials: All datasets used in this article were from publicly available internet sites and their sources are referenced and/or are cited in the article.